OpenStudy (anonymous):

what is the solution to the system of equations? x + y = 5 x - 2y = 2 ( ? , ? )

OpenStudy (aihberkhan):

Okay! I can help! :)

OpenStudy (aihberkhan):

Now lets work with the 1st equation to start off. First we will subtract \(y\) from both sides. When you do that you should get \(x = 5 - y\)

OpenStudy (aihberkhan):

So now, we know what we can substitute \(x\) for in the second equation. So we will plug in \(5 - y\) in for \(x\) in the second equation.

OpenStudy (aihberkhan):

When we do this the equation should look like this: \((5 - y) - 2y = 2\)

OpenStudy (aihberkhan):

Are you following so far? :) @DietSodaa

OpenStudy (anonymous):

yes :)

OpenStudy (aihberkhan):

Awesome! Now lets solve that equation. So we want to isolate\(y\) on one side. So let's subtract 5 from both sides. When you do this our equation should look like: \(-y - 2y = -3\)

OpenStudy (aihberkhan):

So now just combine like terms and finally divide. When do you that you should get \( y= \frac{ 8 }{ 7 }\)

OpenStudy (aihberkhan):

Okay now we know y! :) So we can plug that in for \(y\) in the first equation! :) When you do this the equation should look like: \(x + \frac {8 }{7} = 5\)

OpenStudy (aihberkhan):

Now subtract \(\frac{8}{7}\) from both sides and then just combine like terms. When do you this you should get \(x = \frac{6}{7}\)

OpenStudy (aihberkhan):

So our final answer is: \((\frac{6}{7}. \frac{8}{7})\)

OpenStudy (aihberkhan):

Hope this helped! Have a great day! :) If you see that I am online and need help with a question, just tag me in your question! @DietSodaa